# The Dual Personas of Social Media Bots

### v1

@@ 2 The Personas of the Social Media Bot
We surveyed the literature for the analysis of bot activities to build our taxonomy. We annotated the bot activities in the user-content-interaction frame. Then we derived bot personas that embodied the mechanics employed. These bot personas fall into two big categories: bot personas by behavior and bot personas by content. We characterize each of these personas in terms of the first-principles definitions of a bot, laying out the user-content-interaction characteristics of each bot persona. The fifteen personas and their definitions are reflected in Table 1 and Table 2.
@@
Chaos Bot Sow chaos and discord Higher than average BEND values
@@
Persona Definitions Properties
User Content Interaction
@@
Chaos Bot Sow chaos and discord Higher than average BEND values1
Information Correction Bot Correct information References to fact checking website or presence of negation
Table 2: Personas of social media bots (By Content). 1BEND cues are signals of information maneuver tactics, derived from [1]
@@ 3 The Duality of Social Media Bot
Genre-specific bots in the political realm demonstrate the negative impact of bot agents. Synchronized bots have been deployed to artificially inflate support for political candidates [45, 46], and push specific narratives in a coordinated fashion [47, 48]. Overall, bots manipulate the public discourse [24], and are regarded to have affected election outcomes [49].
@@ 4 Dual Personas of Social Media Bots
Here we discuss the dual personas of the social media bot as observed in literature. These studies are presented in Table 5 and  6. Much of the bot analysis literature focuses on the content of bot agents, observing many content-based bot personas, such as the content generation bot and the genre specific bot. The bot personas found in literature are mostly situated with bad content, such as opinion manipulation [51], inciting online polarization [52] and disinformation spread [8, 23]. In terms of behavior-based bot personas, the amplifier bot [53, 47, 54] and the synchronized bot are most studied [55, 12], particularly in the context of political manipulation and polarization [53, 56].

### v3

@@ 2 The Personas of the Social Media Bot
We surveyed the literature for the analysis of bot activities to build our taxonomy. We annotated the bot activities in the user-content-interaction frame. Then we derived bot personas that embodied the mechanics employed. These bot personas fall into two big categories: bot personas by behavior and bot personas by content. We characterize each of these personas in terms of the first-principles definitions of a bot, laying out the user-content-interaction characteristics of each bot persona. The fifteen personas and their definitions are reflected in Table 1 and LABEL:tab:personas_2. The breakdown of properties of personas served as heuristics for identification of each persona. Such heuristics can also be used for generating the personas, as observed in the AuraSight scenario [41], where simulated social media data were generated.
@@
Chaos Bot Sow chaos and discord Higher than average BEND values1
@@
Finally, we ran these heuristics on a subset of data taken from the 2020-2021 coronavirus vaccine discussions on X. These discussions were based on the conversations that took place the week before the coronavirus vaccine rollout (1 December - 7 Decemeber, 2020). The data was collected through the Twitter V1 Streaming API and filtered for vaccination terms, and consisted of 1,076,734 tweets from 580,135 unique users. We urge the reader to refer to the original article for the data collection details [13]. We applied the heuristics from Table 1 and LABEL:tab:personas_2 on the data, and summarize examples of tweets from each type of bot in Table 2 and Table 3.
Persona Illustrated examples
@@
Chaos Bot From the same user: “DON’T TAKE IT PEOPLE! I spoke to old lady today and her daughter has told her she cant see grand daughter without having vaccine WHAT A XXXXXXX BITCH”, “Former Pfizer Science Officer Reveals Great COVID-19 Scam”
Announcer Bot @GlobalPandemic begins each tweet with “ALERT:”. For example, “ALERT: Novavax, Maryland biotech firm, completes enrollment for COVID-19 vaccine trials […]”, “ALERT: CDC-convened committee votes to add nursing home residents to first phase of COVID-19 vaccine access […]”, “ALERT: U.K. becomes first country to approve Pfizer/BioNTech COVID-19 vaccine […]”
Content Generation Bot @quant_data tweets original tweets about the financial markets “Moderna shares are trading higher on optimism surrounding the company’s COVID-19 vaccine candidate. Related Tickers: \$MRNA [URL]”, “Carnival shares are trading higher as positive COVID-19 vaccine news in recent weeks has boosted investor sentiment and lifted cruise sector recovery hopes. Related Tickers: \$CCL [URL]”
Information Correction Bot @FactCheck tweets: “Terrific questions - thank you. I have put question 1 to several experts and they have told me that they generally expect side-effects from a vaccine injection to show up quickly. Delayed reactions are unusual. (1)”
Genre Specific Bot @IndiaTrendDaily mostly talks about latest gear for sales. Example tweets: “DELL Inspiron 16inch HD Laptop [URL]”, “HP 1.6-inch Laptop [URL]”
Conversational Bot @CuddleTalk: “@greta Greta, anyone taking an mRNA COVID-19 vaccine that modifies one’s mRNA to produce the spike proteins on a COVID-19 […]”, “@GiannoCaldwell Why would anyone take a COVID-19 vaccine that modifies their mRNA to produce the spike proteins on the outside of the COVID-19 cell to produce an immune response?”
Engagement Generation Bot @DonnaYouDC tries to get readers to think about the anti-vaccination arguments: “And if the #pharma #biotech, pharmacy & distributor execs are ”irritated at pressure from the White House to attend an event they perceive to be largely political,” they certainly don’t have to attend. So, will they anyway? \$PFE \$MRNA \$FDX \$UPS \$MCK \$WBA \$CVS #COVID19 #vaccines”, “Why would #pharma #biotech, etc. want Trump to take credit for their work? \$PFE \$MRNA \$FDX \$UPS \$MCK \$WBA \$CVS #COVID19 #vaccines”
News Bot Example users: @dwatchnews_asia, @novasocialnews, @HealthNewsFL
Example tweets: “Pfizer seeks emergency approval for COVID-19 vaccine in India, media say [URL]”; “China prepares large-scale rollout of coronavirus vaccines (from @AP) [URL]”
Table 2: Examples of tweets by personas of social media bots (By Content).
Persona Illustrated examples
Behavior-Based Bot Persona
Social Influence Bot “STOP THE BLOODTHIRSTY LUCIFARIAN GLOBALISTS AND YOUR PLANDEMIC WHICH IS REALLY A…#LucifarianNewWorldOrderTakeover #TheGreatEvilReset #Agenda2030 #Agenda21 #Depopulation #WeDoNotConsentCVVaccine2020 #JesusIsLord #NotBillGates #GodWins”
Amplifier Bot Tweets from the same user @CEPIvaccines: “RT @UN_PGA: We need to join forces to ensure #Vaccines4All. Fair and equitable access to #COVID19 vaccines is both the right thing to do & …”, “RT @gavi: The Gavi #COVAX Advance Market Commitment (AMC) is an innovative financial mechanism designed to ensure that the world’s poorest …”, “RT @gavi: #COVAX is founded on the principle of equitable access because in the #COVID19 vaccine race, nobody wins unless everyone wins. …”
Cyborgs @IAM__Network uses several platforms to tweet (Blog2Social, IAMBLOG2TWITTER, Microsoft Power Platform)
Bridging Bot Example tweets: “@LongJonson2 @JimmyJGreen @wbz If interested I wrote up exactly how these vaccines work. This doesn’t change your DNA at all. It utilizes the ribosomes in your cells to read an mRNA strand to create the protein. This is something our cells do naturally every day https://t.co/fdRTIK7Xie” ; “@RenTradewind @entobird @JUNIUS_64 @ROCKETDRAG @MarziManed @CatsCavern Molecular Biologist here! I study Viral Diseases, Viral Immunity and Pathogenesis for a living at the National Institute of Allergy and Infectious Diseases. Current research efforts have been dedicated to the Coronavirus vaccine. #FurriesinSTEM [URL]”
Repeater Bot @ElsaKekita repeats the same tweet more than 3 times “La vacuna de Moderna otorga al menos 3 meses de inmunidad frente el coronavirus - Infobae [URL]”
Self-Declared Bot Example Users: @CoronaUpdateBot, @Bot_Corona_V, @GothBotAlice
Synchronized Bot @TheNewsGob, @OnlyFilmMedia, @VintageVideoPod, @newsbreakApp frequently tag each other
Table 3: Examples of tweets by personas of social media bots (By Behavior).
@@ 3 The Duality of Social Media Bot
Genre-specific bots in the political realm demonstrate the negative impact of bot agents. Synchronized bots have been deployed to artificially inflate support for polit,ical candidates [46, 47], and push specific tweets from narratives in a coordinated fashion [48, 49]. Overall, bots manipulate the public discourse [24], and are regarded to have affected election outcomes [50].
@@ 4 Dual Personas of Social Media Bots
Here we discuss the dual personas of the social media bot as observed in literature. These studies are presented in Table 6 and  7. Much of the bot analysis literature focuses on the content of bot agents, observing many content-based bot personas, such as the content generation bot and the genre specific bot. The bot personas found in literature are mostly situated with bad content, such as opinion manipulation [52], inciting online polarization [53] and disinformation spread [8, 23]. In terms of behavior-based bot personas, the amplifier bot [54, 48, 55] and the synchronized bot are most studied [56, 12], particularly in the context of political manipulation and polarization [54, 57].